# Photon Number Coherence in Quantum Dot-Cavity Systems can be Enhanced by Phonons

### v1

@@ I Introduction
Quantum dots (QDs) have been shown to be excellent single-photon sources for many applications and have steadily been improved regarding certain figures of merit, such as purity,[1, 2, 3, 4, 5, 6] indistinguishability [6, 7, 8] and brightness.[2, 3, 4, 6] They have, in particular, been used for quantum-cryptographic applications,[9, 10] where the coherence between different photon-number states, known as photon number coherence (PNC) turns out to be a crucial figure of merit.[3] Some protocols, such as the famous BB84 key distribution protocol,[11] benefit from low coherence,[12, 13] while others benefit from the presence of PNC,[3, 14] such as twin-field quantum key distribution with single photons.[15] PNC has so far mostly been investigated using Poisson sources,[12, 13, 16] where usually a separate phase scrambling step is performed,[17, 18] which additionally decreases the secure key rate and can enable unambiguous state discrimination attacks, if it is performed imperfectly.[19, 13] Additionally, Poisson sources must be used at very low brightness to work in the few-photon regime,[20] while QDs provide photons on-demand [21] promising a significantly higher communication rate.[22]
@@ II The QD-cavity system
with a full width at half maximum FWHM=22ln $\text{FWHM}=2\sqrt{2\text{ln}\,2}\,\sigma=3~{}\text{ps}$ of the laser amplitude, and a pulse area $\Theta$, are used. We include the coupling to longitudinal acoustic (LA) phonons, that are known to be the most important ones for typical QDs.[30, 31] The LA phononic environment, where $\hat{b}_{j}$ ($\hat{b}^{\dagger}_{j}$) annihilates (creates) a phonon of mode $j$ with energy $\omega_{j}$, is described by the Hamiltonian [37, 38]
@@
coupling each mode to the exciton with the coupling constant $\gamma_{j}^{X}$.[27, 28] Here, we consider a QD with a size of 3 nm. The phonon parameters of the InGaAs/GaAs QD were chosen as in Ref. [39]. Thus the total Hamiltonian is given by
@@ IV.1 Rabi rotations
Figure 3 shows the integrated one-photon population $\tilde{\rho}_{11}$ along side the PNC $\tilde{\rho}_{01}$ as a function of pulse area $\Theta$ for different temperatures starting from $4.2$ K and then increasing from $10$ K in steps of $20$ K to $110$ K. For temperatures above 8080~{}8080 K optical phonons are expected to play a prominent role,[44] restricting the validity of our model to temperatures $T\lesssim 100$ K. The phonon-free case is displayed in green for reference.
@@
These mechanisms similarly affect the electronic coherence and because of its close link to the photonic coherence, they also affect the PNC, which is shown in Figure 3  b). However, while the electronic coherence is largest for the $\pi/2$-pulse without phonons, it cannot increase when phonons are added. The PNC, in contrast can, since the filter function can have such an effect. As a result, it is possible for $\tilde{\rho}_{01}$ to increase for $\pi/2$ pulses above the phonon-free case. This surprising effect can be traced back to the renormalized coupling strengths,[29, 47, 34, 48] that influence this integrated quantity. Another remarkable effect is seen by looking at the behavior at pulse areas around $\pi$. While the PNC in the phonon-free case reaches almost zero, we find that the PNC remains much larger when phonons are accounted for. This effect is due to the damping of the Rabi rotations. Since a full inversion is not reached, the electronic coherence can still have finite values. At higher pulse areas, the PNC is then strongly damped and no more oscillations are visible, if the temperature is sufficiently large. Thus, depending on the pulse area, phonons can have coherence boosting or damping effects.
@@ IV.2 QD-cavity detuning
Figure 6: a) Dependence of the filter function $F$ of the detuning $\Delta\omega_{CX}$ and the QD-cavity coupling $g$ b)/c) Temporal integration of the photonic 01-number coherence, $\tilde{\rho}_{01}$, displayed for different temperatures $T$ and QD-cavity detunings $\hbar\Delta\omega_{CX}$. The chosen temperatures are 4.2 K as well as 10 K to 110 K in 20 K steps. The phonon-free case is displayed in green. The pulse area is chosen to b) maximize the QD population and to c) minimize the QD polarization after the pulse.
@@ V Conclusions
We have analyzed the PNC as function of pulse area, which for the QD occupation leads to the well-established Rabi rotations. In the phonon-free case, the control of the electronic system is passed on to the PNC. This changes significantly in the case with phonons. Although phonons lead to a damping of the electronic coherence, the PNC can increase compared to the phonon-free case.
Because of the renormalization of the QD-cavity coupling, different temperatures produce qualitatively different behavior of PNC for different QD-cavity detunings. Crucially, for low temperatures, the PNC increases for all values.
Our study shows that the influence of phonons on the PNC is highly non-trivial, as the simple assumption that the PNC exclusively decreases with phonons due to decoherence, does not hold true and other phononic effects must be taken into account. As PNC is crucial for many quantum cryptography applications, it is essential to always consider the influence of the solid-state-environment, which might be detrimental, but also could turn out as helpful.[3]
@@
Acknowledgements
P.C.A.H. is grateful for the advice and guidance he received from Dr. Michael Cossacchi and Dr. Tim Seidelmann.
D.E.R acknowledges financial support from the German Research Foundation DFG through project 428026575 (AEQuDot). We are grateful for fruitful discussions with Y. Karli, F. Kappe, V. Remesh and G. Weihs from University of Innsbruck.

### v2

@@ I Introduction
Quantum dots (QDs) have been shown to be excellent single-photon sources for many applications and have steadily been improved regarding certain figures of merit, such as purity,[1, 2, 3, 4, 5, 6] indistinguishability [6, 7, 8] and brightness.[2, 3, 4, 6] Another interesting characteristic of the emitted photons is the coherence between different photon-number states, known as photon number coherence (PNC). So far, PNC has not been much discussed in the quantum-dot community until it was discovered that PNC is of utmost importance for the performance of many quantum-cryptographic protocols [9, 10, 11]. Some protocols, such as the famous BB84 key distribution protocol,[12] achieve the highest security when PNC is as low as possible,[13, 14] while other protocols require the presence of finite PNC,[3, 9] such as twin-field quantum key distribution with single photons.[15] Thus, the optimal value of PNC depends on the specific cryptographic protocol [3] and how much deviation the protocol can tolerate.[9, 10, 11] PNC has so far mostly been investigated using Poisson sources.[14, 13] In most cases, achieving low PNC is the target. Thus, for Poisson sources usually a separate phase-scrambling step is performed,[16, 17] which additionally decreases the secure key rate and can enable unambiguous state discrimination attacks, if it is performed imperfectly.[18, 14] Additionally, Poisson sources must be used at very low brightness to work in the few-photon regime,[19] while QDs provide photons on-demand [20] promising a significantly higher communication rate.[21]
@@ II The QD-cavity system
with a full width at half maximum FWHM=22ln $\text{FWHM}=2\sqrt{2\text{ln}\,2}\,\sigma=3~{}\text{ps}$ of the laser amplitude, and a pulse area $\Theta$, are used. Since it is embedded into the solid state environment, the quantum dot interacts with phonons, which provide an important dephasing mechanism. This coupling can generally occur due to different coupling mechanisms, in particular Fröhlich coupling to longitudinal optical (LO) phonons, piezoelectric coupling to all acoustic phonons as well as the deformation potential coupling to longitudinal acoustic (LA) phonons.[30] Because the electron and hole distributions in neutral QDs are usually similar, the piezo-electric and Fröhlich couplings are very small and do not produce significant dephasing in GaAs QDs. [26, 27] For these reasons, and since LO phonons are assumed to be off-resonant to the transitions to higher electronic QD states, which would otherwise form a polaron in the QD,[36] we consider only the deformation potential coupling to LA phonons. The LA phononic environment, where $\hat{b}_{j}$ ($\hat{b}^{\dagger}_{j}$) annihilates (creates) a phonon of mode $j$ with energy $\omega_{j}$, is described by the Hamiltonian [37, 38]
@@
coupling each mode to the exciton with the coupling constant $\gamma_{j}^{X}$.[26, 27] It should be noted that Equation (5) is the standard microscopic basis to model phonon-induced processes in QDs, in particular pure dephasing.[26] To calculate the coupling constants, we use a simple model in which it is assumed that the electron and the hole are exposed to a harmonic potential. Then, the carrier wavefunctions are given by Gaussians, which have a characteristic width of $a_{e/h}$ for electron and hole, respectively. The resulting QD-phonon coupling constants are given by [26, 30]
$\gamma_{j}^{X}=\frac{|\boldsymbol{q}_{j}|}{\sqrt{2\hbar\rho\omega_{j}}}\left(D_{e}\,\text{e}^{-\boldsymbol{q}_{j}^{2}a_{e}^{2}/4}-D_{h}\,\text{e}^{-\boldsymbol{q}_{j}^{2}a_{h}^{2}/4}\right),$ (6)
where $D_{e/h}$ are the deformation coupling constants, $\rho$ is the mass density of the material and $\boldsymbol{q}_{j}$ is the wave-vector of the $j$th phonon mode, which has a frequency of $\omega_{j}$. This model has given excellent agreement with previous experiments with high sensitivity to the QD-phonon coupling.[39] We chose $a_{e}=3~{}\text{nm}$, which corresponds to a FWHM of the electron density of 7 nm. Here, we assume the same harmonic potential for holes, resulting in $a_{h}=\frac{m_{h}}{m_{e}}\,a_{e}\approx 1.15\,a_{e}$. It has been found that it is also possible to describe non-spherical QDs using this formula when $a_{e/h}$ are appropriately chosen.[40] All other phonon parameters of the InGaAs/GaAs QD were chosen as in Ref. [41]. Thus the total Hamiltonian is given by
@@
In order to obtain specific predictions starting with a Hamiltonian of the type Equation (7) numerous methods are in use, like, e.g., non-equilibrium Green’s function [44], the correlation expansion of density matrices [45, 46, 47, 48], time-convolutionless projection operators [49] or the polaron master equation [50, 35] to name but a few. Most of these methods introduce approximations to the model to make the numerics feasible. For example, the correlation expansion accounts only incompletely for multi-phonon processes which leads, e.g., to missing structures in corresponding spectra [26]. The correlation expansion also has limits when strong phonon coupling or long simulation times are considered [51]. The polaron master equation involves a Markov approximation in the polaron transformed frame which makes this approach unreliable at strong time-dependent driving. As a result, the standard polaron master equation does not correctly capture, e.g. the reappearance of Rabi rotations [35] as predicted in Ref. [52] that has recently been measured [53].
Indeed, it is challenging to account for the wealth of phenomena induced by the QD-phonon coupling that go far beyond simple pure-dephasing obtained by weak-coupling master equations. This includes phonon-assisted transitions and thermalization, polaron formation, which dresses and renormalizes driving and coupling terms in the QD Hamiltonian, and non-Markovian quantum dynamics. Moreover, also the effective dephasing rate strongly depends on details of the driving, which is particularly challenging to work with in the case of pulsed driving.
For simple systems, such as a two-level system coupled to an oscillator bath, the pioneering work of Makri and Makarov [54, 55] has shown that a numerical treatment without approximation to the model is possible by representing the time-evolution operator of the system formally exact as a path integral and performing the necessary sum over the paths using an iterative scheme. For the model considered in the present paper, however, the original algorithm would require the iteration of estimated $\sim 10^{11}$ complex valued numbers. Thus, we use for our numerics a modified algorithm [56] that reduces in our case the number of elements that need to be iterated by about 5 orders of magnitude. Our modified path-integral approach can be formulated in Liouville-space [41], as needed when Markovian rates e.g. for cavity losses and radiative decay should be accounted for, as we do in Equation (9). It is also suitable for calculating multi-time correlation functions [57], which is necessary for the calculation of the purity and indistinguishability. Our implementation of path-integrals for the simulation of multi-time functions avoids the use of the quantum regression theorem (QRT). The QRT involves a Markov approximation even when the memory for the single time propagations is fully accounted for, which can introduce sizable errors.[31, 58] We are able to obtain converged results that invoke no approximation to the model such that only well controlled discretization errors occur.
@@ IV.1 Rabi rotations
Figure 3 shows the integrated one-photon population $\tilde{\rho}_{11}$ along side the PNC $\tilde{\rho}_{01}$ as a function of pulse area $\Theta$ for different temperatures starting from $4.2$ K and then increasing from $10$ K in steps of $20$ K to $110$ K. For temperatures above8080~{}8080 K, optical phonons are expected to play a prominent role,[59] restricting the validity of our model to temperatures $T\lesssim 100$ K. The phonon-free case is displayed in green for reference.
@@
These mechanisms similarly affect the electronic coherence and because of its close link to the photonic coherence, they also affect the PNC, which is shown in Figure 3  b).
First, we consider $\pi$-pulses, where the PNC reaches almost zero in the phonon-free case. We find that the PNC remains much larger when phonons are accounted for. This effect is due to the damping of the Rabi rotations. Since a full inversion is not reached, the electronic coherence can still have significant values. And because the electronic coherence is the source of the PNC, the latter also increases.
When we instead consider $\pi/2$-pulses, where without phonons PNC is maximal, we observe that even here an increase of PNC is possible. This surprising effect can be traced back to the renormalized coupling strengths,[28, 33, 62, 63] that influence this integrated quantity. The reason for this is that $F$ in Equation (16) increases for decreasing coupling and the coupling decreases due to the renormalization effect of phonons. Physically this is due to the slow-down in the emission of the coherence out of the QD-cavity system. Note that this effect of course plays a role for all pulse-areas, since $F$ does not depend on the initial excitation. This effect therefore also influences the increased value of PNC for $\pi$-pulses.
At higher pulse areas, the PNC is then strongly damped and no more oscillations are visible, if the temperature is sufficiently large. Thus, depending on the pulse area, phonons can have coherence boosting or damping effects.
@@ IV.2 QD-cavity detuning
Figure 6: a) Dependence of the filter function $F$ of the detuning $\Delta\omega_{CX}$ and the QD-cavity coupling $g$ b)/c) Temporal integration of the photonic 01-number coherence, $\tilde{\rho}_{01}$, displayed for different temperatures $T$ and QD-cavity detunings $\hbar\Delta\omega_{CX}$. The chosen temperatures are 4.2 K as well as 10 K to 110 K in 20 K steps. The phonon-free case is displayed in green.
The pulse area is chosen to b) maximize the QD population and to c) minimize the QD polarization after the pulse.
@@
[…]
These results are encouraging despite the phonon-induced increase in PNC at low temperatures, when keeping other figures of merit of single-photon sources, such as purity, indistinguishability and brightness, in mind. In most cases, one is interested in having low PNC. While the minimally achievable PNC decreases for larger temperatures, the effect is not very pronounced in Figure 4 and one can reasonably use low temperatures without too large a penalty to PNC. In this case, very good values for purity, indistinguishability and brightness are possible [31, 64] because the $\pi$-pulse area is approximately equal to the pulse area minimizing PNC for low temperatures. When instead, PNC should be maximized, one needs to use a pulse area similar to $\pi/2$. In this case, purity and indistinguishability still have high values, but the brightness suffers a significant penalty of about $1/2$.[64] This reduction, however, must be accepted when electronic coherences are required e.g. for coherent control.
In order to quantify how important figures of merit perform for the parameters optimized for low or high PNC values, we have calculated purity and indistinguishability for the same parameters as in Figure 4 for $T=4.2\text{ K}$. Details how these quantities are calculated are given in Ref. [31]. As seen in Figure 7, the purity is not significantly affected by changes in the QD-cavity detuning and stays near unity regardless whether PNC is optimized for high or low values. Its values are about $98\%$, which is typical for similar systems[31, 65]. In particular, the results for $\pi$-pulses are almost the same as in the case minimizing PNC, which is likely due to the very similar pulse area needed for these cases. The pulse area maximizing PNC, which is roughly similar to the $\pi/2$ case, has higher purity because for lower driving strengths, the reexcitation probability is reduced. The indistinguishability reaches its highest values for QD-cavity detuning $\hbar\Delta\omega_{CX}=0\text{ meV}$ and falls off for larger detunings in all considered cases. The indistinguishability values in Figure 7 are in reasonable agreement with other results reported for similar systems.[31, 65, 66] For cavity-based systems it has been reported in Ref. [65] that almost perfect indistinguishabilities and purities can be achieved by using the Swing-UP of quantum EmitteR (SUPER) [67] scheme for the preparation instead of Gaussian pulses.
These results demonstrate that for the parameters where the lowest PNC is reached, it is usually possible to simultaneously obtain high values for purity, indistinguishability and brightness. Also, when high PNC is the target, the achievable purity is generally rather high for low temperatures, while the indistinguishability has reasonable values that still depend on parameters like the QD-cavity detuning and details of the excitation. And naturally, for pulse areas near $\pi/2$ the brightness is significantly lower than for $\pi$-pulses.
@@ V Conclusions
We have analyzed the PNC as function of pulse area, which for the QD occupation leads to the well-established Rabi rotations. In the phonon-free case, the control of the electronic system is passed on to the PNC. This changes significantly in the case with phonons. Dispite the well-known pure dephasing effects, phonons were found to sometimes increase PNC. For $\pi$-pulses the reason is the damping of Rabi rotations and for $\pi/2$-pulses it is mostly due to the renormalization of QD-cavity coupling, which slows the emission of the coherence.
Because of the renormalization of the QD-cavity coupling, different temperatures produce qualitatively different behavior of PNC for different QD-cavity detunings. Crucially, for low temperatures, the PNC increases for all pulse areas.
Our study shows that the influence of phonons on the PNC is highly non-trivial, as the simple assumption that the PNC exclusively decreases with phonons due to decoherence, does not hold true and other phononic effects must be taken into account. As PNC is crucial for many quantum-cryptography applications, it is essential to always consider the influence of the solid-state-environment, which might be detrimental, but also could turn out as helpful.[3]